# Continual Learning of Large Language Models: A Comprehensive Survey

### v1

@@ 4.1.3 Distributional Shifts in CPT
Medical HuatuoGPT-II [42] DAP + U-SFT Baichuan2 Replay ZS ZS | HE
Financial XuanYuan 2.0 [355] DAP + SFT BLOOM Replay HE HE
Scientific PLlama [331] DAP $\rightarrow$ IT GAL Replay L L | ZS
E-Commerce EcomGPT-CT [187] DAP $\rightarrow$ SFT BLOOM Replay ZS | FS ZS | FS
Legal Layer Llama [115] DAP $\rightarrow$ G-IT $\rightarrow$ D-IT Llama Replay ZS ZS
Multi-Domain AdaptLLM [49] DAP Llama Replay ZS ZS | FT
4.2 Domain-Adaptive Pre-training (DAP)
Background of DAP. Institutions, regardless of size, often possess significant amounts of unlabeled, domain-specific data. This data bridges the gap between general-purpose LLMs trained on diverse corpora and fine-tuned LLMs designed for specific downstream tasks. Leveraging this data as a preparatory stage can facilitate effective adaptation of LLMs to downstream tasks. Such process of “continued/continual/continuous pre-training” [327, 88, 187, 93, 323, 319, 115, 177, 320, 10, 344, 52, 350], “further pre-training” [269, 166, 63, 246, 3], “domain tuning” [243], “knowledge enhancement pre-training” [177], and “knowledge injection training” [311] is unified and termed “Domain Adaptive Pre-training (DAP)” [92] for clarity and consistency throughout the remainder of this survey. In the pioneering work of domain-adaptive pre-training (DAPT) [92], the authors continuously pre-train the language models on a larger domain-specific dataset before fine-tuning them to the downstream tasks, resulting in universally improved performance aross various tasks. As the observation above has been validated on multiple domains in parallel, including BioMed, CS, News, and Reviews [92], practitioners commonly accept that employing DAP on additional unlabeled domain-specific data benefits downstream tasks. Consequently, this technique has become widely deployed in many modern LLMs.
Summary of LLMs with DAP. To illustrate this, we provide a summary of existing studies utilizing domain-adaptive pre-training for LLMs in Table 2. Each entry in the table is characterized by three main features: (i) training process specifications, encompassing the vertical domain for which LLMs are trained, the training pipeline preceding release, and the LLM architecture employed; (ii) adopted continual learning techniques, including rehearsal, parameter regularization, and architecture expansion; and (iii) evaluation metrics for continual learning, such as backward transfer (forgetting) and forward transfer (adaptation to downstream data). Following this overview, we will present general observations about DAP in Section 4.2.1, followed by a detailed introduction to LLMs developed in vertical domains in Section 4.2.2.
4.2.1 General Observation on DAP
As depicted in Table 2, several key observations emerge regarding the current research landscape of DAP. Firstly, DAP predominantly occurs in a single stage. Horizontal Continual DAP which involves more than one stage is seldom explored: among the 34 papers listed, only one paper employs two stages of DAP [245]. In Code Llama [245], aimed at developing a language model tailored to Python programming, the authors initialize the model from the pre-trained Llama 2 checkpoint. They then conduct the first stage of DAP across multiple programming languages (500 billion tokens) before proceeding to the second stage, focusing solely on Python code (100 billion tokens). Finally, they perform long context fine-tuning (20 billion tokens) to enhance the model’s capability in challenging long-context scenarios of code generation. This PT $\rightarrow$ DAP $\rightarrow$ DAP $\rightarrow$ FT pipeline represents the sole example found thus far that strictly adheres to the definition and hierarchical structure of vertical continuity in pre-training and adapting LLMs for final end-use. Hence, categorizing the 10 studies that solely conduct one stage of DAP and nothing more [177, 209, 269, 323, 158, 87, 326, 221, 10, 49] proves challenging. One could also argue that they deploy an additional single stage of CPT rather than DAP. Nevertheless, considering that all these papers aim to adapt a general-purpose LLM to a specific domain, we include them in this section for discussion, aligning with the categorization we have established thus far.
Secondly, the notion of interpreting DAP through the lens of continual learning, whether intentional or not, is widely embraced. As demonstrated in Table 2, with the exception of the first section (white, 11/33), where papers overlook any potential side effects of DAP leading to vertical forgetting of previously learned general knowledge, the remaining sections (all gray, 22/33) either evaluate the potential negative impacts of DAP or proactively employ continual learning techniques to mitigate the risk of vertical forgetting from the outset.
Thirdly, we observe widespread adoption of CL techniques (14/33) for training domain-specific LLMs. However, the diversity of these techniques is limited, with only replay [52, 311, 10, 243, 42, 355, 331, 187, 115, 49] and parameter expansion (LoRA [326, 221, 310, 327]) or Layer/Block expansion [310, 327] being utilized. This highlights the need for further research to investigate, incorporate, and design more sophisticated CL techniques for not just DAP, but vertical continual learning in general. In fact, it appears that individuals may not explicitly recognize that DAP should be viewed from the perspective of vertical continuity, as they often employ CL techniques unknowingly. This deduction arises from two observations: (i) parameter expansion methods inherently embody implicit CL techniques. For instance, in LoRA [110], the increment of weights $\Delta{\bm{W}}={\bm{O}}$ preserves the original performance on previous data distributions, but once adaptation occurs ($\Delta{\bm{W}}\neq{\bm{O}}$), forgetting on the original data distribution follows. This analysis extends to other parameter expansion techniques such as layer expansion [327] and block expansion [310]. Authors typically empirically demonstrate the effectiveness of these approaches, attributing forgetting mitigation to the low-rank property and parameter efficiency; (ii) excluding parameter expansion methods, replay emerges as the only CL technique employed during DAP, except in cases where extensive empirical investigations of CL methods are conducted [243]. Furthermore, studies deploying replay often term the technique as “data combination” [311] or “data mixing/mixture” [10, 331, 187, 49], without recognizing it as a vertical continual learning problem.
4.2.2 Different Domains of DAP
Legal Domain. Given the legal industry’s demand for managing ever-growing volumes of legal documents, there’s a burgeoning need to harness LLMs to aid legal professionals in navigating, interpreting, and generating high-quality legal materials [318, 251, 343]. While general-purpose LLMs may perform adequately on some legal benchmarks [191], customizing LLMs with additional unlabeled resources specific to the legal domain can yield superior results. This is because the high-volume unlabeled legal corpus resembles the conditions under which general-purpose LLMs are pre-trained. In Layer Llama [115], the authors gathered publicly available legal texts from China Courts websites, including judgment documents, legal articles, judicial interpretations, court news, and law popularization articles, totaling approximately 10 billion tokens as noted in a GitHub issue [116]. In SaulLM [52], the authors collected the DAP corpus from various jurisdictions in different countries, such as the U.S., Europe, and Australia, resulting in a corpus of 30 billion tokens to cover diverse aspects of legal texts. When combined with previously available datasets [^3] Detailed web links for data collection in SaulLM are omitted here for brevity. Refer to the original paper for more information. [79, 141], the total tokens used for legal-domain DAP reach 94 billion.
The substantial volume of DAP data, while offering valuable insights into specific domains, increases the risk of catastrophic forgetting of the general knowledge due to the large number of update steps involved. To mitigate this, SaulLM incorporates general data from Wikipedia, StackExchange, and GitHub into the DAP data, constituting about 2% of the final dataset [52]. Following DAP, SaulLM then employs a combination of general-domain and legal-domain instructions to enhance the model’s instruction-following ability (U-IT, see Table 2). Similarly, Lawyer Llama also incorporates general-domain data during DAP, but the replay rate is not disclosed [115]. After DAP, it undergoes two distinct phases of instruction tuning: first, general-domain instruction tuning (G-IT), followed by domain-specific downstream legal consultation application instruction tuning (D-IT). However, no explanation is provided for why two-stage IT is preferred over consolidation into one as in SaulLM [52], leaving this as an open question for future research.
Medical Domain. The development of LLMs holds promise for revolutionary changes in the medical industry, offering potential improvements in efficiency and quality across medical communication, disease diagnosis, and decision-making for doctors. While some instances of general-domain LLMs have shown success in providing useful advice and accelerating diagnosis progress for patients [160, 262], direct deployment poses risks. These risks include the potential for sub-optimal solutions, such as imprecise medical advice, and the possibility of harm, such as incorrect drug recommendations and the propagation of medical misinformation [121, 45]. Efforts have been made to develop medical specialists by either collecting medical-domain data and training an LLM from scratch [94, 263, 86, 182], or fine-tuning publicly available LLMs to meet specific medical needs [185, 311, 42, 324, 16, 354]. Among these approaches, domain-adaptive pre-training techniques have been extensively utilized to preserve the communication and instruction-following abilities of a general LLM, preparing it for subsequent medical applications [185, 311, 42].
BioMedGPT [185] is a multi-modal biomedical language model that integrates representations of human language and the language of life (molecules, proteins, cells, genes, etc.). Prior to final multi-modal supervised fine-tuning, the authors initialize the model from Llama2-Chat [280] and conduct DAP using extensive biomedical documents from S2ORC [172], without considering any continual learning techniques or metrics. In [88], a Chinese medical LLM is developed for medical question answering. DAP is performed using Chinese medical encyclopedias and online expert articles, yielding over 364,000 question-answer pairs, with next-token prediction as the sole training objective. Results show that for both models initialized from Llama-2 [280] and Chinese-Llama-2 [345], performance gradually deteriorates on general-domain datasets as training steps increase, while improving on the Chinese medical examination evaluation [105, 169, 151]. To ensure the model possesses sufficient fundamental medical knowledge before the instruction tuning phase, PMC-LLama [311] gathers biomedical papers from S2ORC [172] and medical textbooks for “knowledge injection training”. During this phase, a general language corpus from RedPajama-Data [53] is replayed at a 5% rate within a training batch. However, the paper does not analyze the effectiveness of adding and mixing general-domain data during DAP.
To mitigate vertical forgetting, AF Adapter [327] proposes an adapter structure extending the original transformer block. This structure extends the width of Attention layers and FFNs for domain-specific knowledge storage. During DAP, the general-domain parameters inherited from the BERT checkpoint remain frozen, while only the adapters are trained. To gauge the extent of forgetting after DAP, the authors collect a small subset of general-domain samples from WuDao [341] and calculate the accuracy of masked word prediction using the same pre-training protocol as BERT and DAP of AF Adapter. They find minimal performance degradation compared to other full-parameter fine-tuning techniques. HuatuoGPT-II [42], proposes to fuse the DAP into the final SFT, making the two-stage development one unified protocol. The challenge of such process mainly comes from the data heterogeneity of the domain-adaptive pre-training unlabeled corpus. The authors address this challenge by reformulating paragraphs of data into (instruction, output) format using existing large language models. During the unified one-stage SFT, they employ a priority sampling strategy to avoid compromising downstream ability, as seen in fixed-rate data mixing [280]. Unlike previous approaches, this paper empirically demonstrates the superiority of unified one-stage SFT over two-stage training, questioning the current computational paradigm of adapting LLMs to specific domains via DAP. In [243], the authors investigate the effectiveness of continual learning techniques for mitigating catastrophic forgetting in DAP on medical-domain corpus. They find that LMs constrained by continual learning techniques on source domains exhibit greater robustness to future domain shifts. Specifically, they identify that parameter regularization techniques like EWC [140], despite slightly higher computational cost, can facilitate beneficial forward and backward transfer, offering valuable insights to the community.
@@ 4.2.2 Different Domains of DAP
Financial Domain. Similar to the medical domain, large language models hold immense potential for enhancing financial communication, decision-making processes, and risk assessment for both traders and ordinary individuals [256, 333, 289, 152]. However, the financial domain entails high stakes, with even minor errors bearing significant consequences. Thus, integrating LLMs into financial workflows requires heightened caution to avoid inaccuracies or misunderstandings that could lead to substantial financial losses or the dissemination of misleading information. Despite advancements, a gap persists between general-purpose large language models and existing domain-specific smaller-scale language models [8, 312], underscoring the urgent need for more powerful financial-domain experts through the integration of LLMs. Notably, DAP techniques have emerged as crucial tools for tailoring LLMs to the intricacies of the financial domain while mitigating the negative effects of abrupt domain shifts from general to finance [177, 152, 323, 326, 355].
@@
BBT-Fin [177] collects a Chinese financial DAP dataset comprising 80 billion tokens sourced from corporate reports, analyst reports, social media, and financial news. In addition to the conventional masked language modeling (MLM) training objective, BBT-Fin further incorporates triplet masking and span masking techniques during DAP. This knowledge enhancement pre-training entails: (i) actively selecting sentences containing specific knowledge triplets (head entity, relation, tail entity), masking one of them, and (ii) simultaneously masking 15% of a random-length span. CFGPT [152] creates CFData, a financial dataset for domain-adaptive pre-training and supervised fine-tuning, comprising 141 billion tokens. The dataset includes corporate prospectuses, announcements, research reports, social media content, financial news, and Wikipedia articles. During DAP, CFGPT does not employ continual learning techniques but utilizes QLoRA [66] for preventing overfitting to downstream data and balancing general response ability and domain-specific ability during supervised fine-tuning. These two methods are typical domain-specific LLM studies focusing solely on target domains, without explicit continual learning measures or evaluation of performance degradation on general skills.
In [323], the authors aim to enhance the data efficiency of domain-adaptive pre-training. They propose two data selection techniques: (i) efficient task-similar (ETS) and (ii) efficient task-agnostic (ETA) domain-adaptive pre-training. These methods are inspired by the generalization bound of domain adaptation [18, 77, 260]. Suppose the source domain data distribution is ${\mathcal{D}}$ and the task data distribution is ${\mathcal{T}}$. The generalization error $\epsilon_{{\mathcal{T}}}(h)$ of the hypothesis $h$ on the target task distribution ${\mathcal{T}}$ can be given as
$\epsilon_{{\mathcal{T}}}(h)$ $\leq\epsilon_{{\mathcal{D}}}(h)+\tfrac{1}{2}d_{{\mathcal{H}}\Delta{\mathcal{H}}}({\mathcal{D}},{\mathcal{T}})+{\mathcal{C}},$ (13)
@@
where $\epsilon$ is the 0-1 error function for binary classification, $h$ represents the model, $d_{{\mathcal{H}}\Delta{\mathcal{H}}}$ is the ${\mathcal{H}}\Delta{\mathcal{H}}$ divergence that measures the distributional discrepancy between two distributions based on a hypothesis set ${\mathcal{H}}$ used for discriminating different data distributions [18], and ${\mathcal{C}}$ is a constant. Therefore, based on the theory, finding a set of examples from the DAP corpus that are similar to the downstream task’s data, i.e., $d_{{\mathcal{H}}\Delta{\mathcal{H}}}({\mathcal{D}},{\mathcal{T}})$ is low, should help the final-stage adaptation to the tasks. In the context of large language models (LLMs), [323] suggests ensuring novelty and diversity in the sampled corpus for DAP. Here, high perplexity of an LLM on a sentence indicates high novelty, while high entropy of part-of-speech (POS) tags on a sentence indicates high diversity [19, 348]. This approach significantly enhances DAP efficiency: it utilizes only 10% of the originally collected data yet outperforms models trained on the entire dataset, underscoring the importance of data quality over quantity in DAP. Furthermore, with the reduced number of examples, the authors do not observe any signs of forgetting regarding the model’s capacity in open-domain scenarios. WeaverBird [326] introduces an intelligent finance dialogue system, where the encoder is trained on Chinese and English financial documents, alongside expert-annotated financial query-response pairs, using LoRA [110]. Xuanyuan 2.0 [355], akin to HuatuoGPT-II [42], proposes the technique of hybrid-tuning, which fuses the stage of DAP and SFT into one, general-domain data and financial-domain data into one. Notably, the distribution of data in hybrid-tuning is non-conventional: financial pre-training data comprises 13%, general pre-training data 20%, and instruction tuning data 67%. Within the instruction tuning data, the financial domain accounts for approximately 30%. This prompts a pertinent question in line with the investigation on efficient DAP in [323]: Is a large DAP dataset necessary for developing a domain-specific LLM?
Scientific Domain. Vertical scientific large language models [276, 336, 322, 350] span many subjects, including astronomy [209, 224], mathematics [10, 338, 181, 344, 84], geology [78, 242, 166, 286], chemistry and physics [246], biology [331, 360, 32, 1, 40, 20, 367]. However, among all the studies listed above, only a small fraction of them adopt the technique of DAP.
@@
OceanGPT [20] is the first LLM tailored specifically for the ocean domain, catering to a range of downstream ocean science applications. It compiles a raw corpus of ocean science literature, prioritizing recent research and historically significant works, and performs DAP on the Llama 2 model. K2 [63] pioneers the development of a foundational language model tailored specifically for geoscience. It aggregates geoscience open access literature and Earth science-related Wikipedia pages for DAP. Following this, it undergoes multi-task instruction tuning utilizing LoRA [110] on both a general instruction tuning dataset and the GeoSignal benchmark introduced within the K2 framework. AstroLlama [209] gathers abstracts solely from astronomy papers on arXiv and proceeds with pre-training focused on next token prediction. It observes an improved perplexity on the domain of scholarly astronomy, without providing more quantitative evaluation. MarineGPT [367] is a multi-modal LLM designed specifically for the marine domain. During DAP, MarineGPT incorporates 5 million marine image-text pairs to imbue domain knowledge. This involves training the parameters of a Q-Former [153] between the frozen visual encoder [71] and text decoder [279]. The four methods introduced above lack further validation regarding the necessity of DAP deployment. They seem to adopt DAP merely out of convention, likely for the convenience of utilizing domain-specific corpora as pre-training data.
@@
Another branch of methods proactively integrate in the replay of the general-domain data to mitigate vertical forgetting. GeoGalactica [166] introduces a series of large language models tailored for geoscience. In the DAP phase, besides the 52-billion-token geoscience corpus, Arxiv papers and Codedata are incorporated, with a mixing ratio of 8:1:1. The authors believe that the inclusion of the Codedata during the model’s pre-training can significantly boost the reasoning ability of the LLMs, which is crucial to fine-tuning on the downstream tasks. However, although GeoGalactica pinpoints challenges of DAP, including overfitting, catastrophic forgetting, maintaining the training stability and convergence speed, it does not further provide empirical evidence supporting the inclusion of the Codedata, or deploying specific measures to address the challenges proposed above. Llemma [10] focuses on mathematics, initialized from Code Llama [245], and undergoes DAP on a blend of the 55 billion mathematical pre-training dataset (Proof-Pile-2 including scientific papers, web data, and mathematical code) and general domain data (Pile [79]) at a 19:1 ratio. In contrast, PLlama [331], designed for plant science, mixes domain-specific and general-domain data at a ratio of 9:1.
@@
Code Domain. The development of LLMs for automatic code filling, debugging, and generation holds significant practical importance, with programmers worldwide benefitting from recent advancements in coding assistants [206, 272]. These advancements cover various frameworks, including encoder-only [206], encoder-decoder [296, 293, 35], and decoder-only [215, 214, 366, 43, 158, 176, 87]. As noted in [272], in the era of LLMs, there’s a growing trend towards decoder-only architectures, leveraging models pre-trained on general natural language like Llama [279, 280]. Additionally, there’s a shift in the training objective from utilizing code structures to simpler tasks like next token prediction and infilling.
@@
From a continual learning perspective, the code domain presents unique advantages and challenges for domain-adaptive pre-trainig (DAP) compared to other domains discussed so far. On one hand, its hierarchical structure (general domain corpus $\rightarrow$ multi-language code $\rightarrow$ specific programming language) provides an ideal training pipeline for DAPs [245], offering potential for more efficient training strategies. On the other hand, programming languages adhere to strict grammars, unlike the more fuzzy and context-dependent nature of natural language. Consequently, language models should ideally leverage these structures through tailored designs. However, given the vast availability of code for training, particularly in popular languages like Python and Java, adopting the same training objectives as for natural languages may yield sub-optimal results. Therefore, many existing works in the field omit domain-adaptive pre-training [296, 293, 186, 207, 123, 305, 372, 69, 155]. In the following section, we will introduce existing code LLMs that employ DAP before the final downstream tasks, discussing both their common attributes and unique characteristics.
@@
Representing a series of notable works that focus solely on adaptation to target domains, CodeGen [215] comprises a suite of LLMs designed for natural language (CodeGen-NL), multi-lingual programming languages (CodeGen-Multi), and mono-lingual programming languages (CodeGen-Mono). These models are trained sequentially, with each subsequent model initialized from the previous one trained on more general-domain data. Specifically, CodeGen-NL is trained on Pile [79], while CodeGen-Multi further trains on a subset of the BigQuery dataset [^4] Link to BigQuery: https://cloud.google.com/bigquery/public-data., which includes source code from six popular programming languages (C, C++, Go, Java, JavaScript, and Python). Subsequently, CodeGen-Mono is trained on BigPython, a large-scale corpus of Python code. Throughout the training process, consistent pre-training objectives are employed, focusing on next token prediction. Comment-Aug [269] addresses the challenge of aligning programming languages with natural languages (PL-NL alignment). Recognizing that denser comments in existing code could enhance the model’s generation capabilities, Comment-Aug proposes a self-augmentation strategy. Initially, it enhances LLMs with comment generation ability through instruction tuning. Then it augments sparsely commented code by generating additional comments, followed by further pre-training on the enriched code data. StarCoder [158] introduces two models: StarCoderBase and StarCoder. StarCoderBase is initially trained on a mixed dataset comprising various programming languages. With the aim of benefiting a broader user base, it maintains the original data distribution without significant reweighting. Subsequently, StarCoderBase undergoes further fine-tuning on an additional 35 billion tokens of Python code, resulting in the development of StarCoder. DeepSeek-Coder-v1.5 [87] originates from DeepSeek-LLM [62] and undergoes pre-training on 2 trillion tokens, comprising 87% source code, 10% English code-related natural language, and 3% Chinese natural language corpus. Unlike Deepseek-Coder, which employs both next token prediction and fill-in-the-middle objectives, DeepSeek-Coder-v1.5 focuses solely on next token prediction during the stage of DAP. Despite this change, initialization from a general-domain LLM results in improved performance across various tasks, including natural language and mathematical reasoning, with minimal performance degradation on coding tasks. This underscores the efficacy of DAP before final downstream tasks in achieving balanced performance across general and domain-specific domains.
As the only work investigated so far that utilizes the general data replay to mitigate vertical forgetting in the code domain, Code Llama [245] introduces a sophisticated training framework tailored for various coding tasks and model sizes, including 7B, 13B, 34B, and 70B variants. Initialized from Llama 2 weights, these models undergo continuous pre-training on a dataset composed of deduplicated public code, discussions about code, and a subset of natural language data. This mix of natural language data serves as a form of pseudo-replay to maintain the models’ proficiency in understanding natural language. During DAP, Code Llama optimizes two objectives: autoregressive next token prediction and code infilling prediction (except for the 34B model, which excludes infilling). Subsequently, Code Llama undergoes further refinement through long-context fine-tuning (LC-FT) to enhance its repository-level understanding. Building upon Code Llama, Code Llama-Instruct undergoes additional instruction-tuning, while a Python-specific variant of Code Llama undergoes additional domain-adaptive pre-training on Python language datasets before LC-FT.

### v2

@@ 4.2.2. Different Domains of DAP
Medical Hippocrates (acikgoz2024hippocrates, ) DAP $\rightarrow$ IT $\rightarrow$ MA Llama2 Mistral LoRA ZS | FS
Language Sailor (dou2024sailor, ) DAP Qwen1.5 Replay ZS
[…]
Medical HuatuoGPT-II (Chen2023HuatuoGPTII, ) DAP + U-SFT Baichuan2 Replay ZS ZS | HE
Financial XuanYuan 2.0 (Zhang2023xuanyuan, ) DAP + SFT BLOOM Replay HE HE
Scientific PLlama (Yang2023PLLaMa, ) DAP $\rightarrow$ IT GAL Replay L L | ZS
E-Commerce EcomGPT-CT (ma2023ecomgptct, ) DAP $\rightarrow$ SFT BLOOM Replay ZS | FS ZS | FS
Legal Layer Llama (huang2023lawyer, ) DAP $\rightarrow$ G-IT $\rightarrow$ D-IT Llama Replay ZS ZS
Multi-Domain AdaptLLM (cheng2024adapting, ) DAP Llama Replay ZS ZS | FT
Language Swallow (fujii2024continual, ) DAP Llama2 Replay FS FS
Financial (takahashi2024pretraining, ) DAP Llama2 Replay Loss | ZS Loss | ZS | FS | RAG
Medical Me-Llama (xie2024me, ) DAP $\rightarrow$ IT Llama2 Replay ZS | FS ZS | FS | FT
Language Aurora-M (nakamura2024aurora, ) DAP $\rightarrow$ IT StarCoder Replay ZS ZS | FS | HE
@@
Financial Domain. Similar to the medical domain, LLMs hold immense potential for enhancing financial communication, decision-making processes, and risk assessment for both traders and ordinary individuals (shah2023zero, ; Yang2023InvestLM, ; Wang2023FinGPT, ; li2023cfgpt, ). Despite advancements, a gap persists between general-purpose LLMs and existing domain-specific smaller-scale LLMs (araci2019finbert, ; Wu2023BloombergGPT, ), underscoring the urgent need for more powerful financial-domain experts through the integration of LLMs. Notably, DAP techniques have emerged as crucial tools for tailoring LLMs to the intricacies of the financial domain while mitigating the negative effects of abrupt domain shifts from general to finance (Lu2023BBTFin, ; li2023cfgpt, ; xie2023efficient, ; Xue2023WeaverBird, ; Zhang2023xuanyuan, ).
@@
BBT-Fin (Lu2023BBTFin, ) collects a Chinese financial DAP dataset comprising 80 billion tokens sourced from corporate reports, analyst reports, social media, and financial news. In addition to the conventional masked language modeling (MLM) training objective, BBT-Fin further incorporates triplet masking and span masking techniques during DAP. CFGPT (li2023cfgpt, ) creates CFData, a financial dataset for DAP and SFT, comprising 141 billion tokens. During DAP, CFGPT does not employ CL techniques but utilizes QLoRA (dettmers2023qlora, ) for preventing overfitting to downstream data and balancing general response ability and domain-specific ability during SFT. These two methods are typical domain-specific LLMs focusing solely on adaptation to target domains without explicit CL measures or evaluation of vertical forgetting.
@@
In (xie2023efficient, ), the authors aim to enhance the data efficiency of DAP. When the downstream tasks’ data distribution ${\mathcal{T}}$ are known, based on the generalization bound (ben2010theory, ; ganin2016domain, ; shi2024unified, ), the authors propose to sample the subset of DAP data whose distribution ${\mathcal{D}}$ is similar to the downstream task’s data, i.e., $d_{{\mathcal{H}}\Delta{\mathcal{H}}}({\mathcal{D}},{\mathcal{T}})$ is low. When the downstream data distribution is unknown, the authors suggest ensuring novelty and diversity in the sampled corpus for DAP. This approach significantly enhances DAP efficiency: it utilizes only 10% of the originally collected data yet outperforms models trained on the entire DAP dataset, underscoring the importance of data quality over quantity. WeaverBird (Xue2023WeaverBird, ) introduces an intelligent finance dialogue system, where the encoder is trained on Chinese and English financial documents, alongside expert-annotated financial query-response pairs, using LoRA (hu2022lora, ). Xuanyuan 2.0 (Zhang2023xuanyuan, ), akin to HuatuoGPT-II (Chen2023HuatuoGPTII, ), proposes the technique of hybrid-tuning, which fuses the stage of DAP and SFT into one, general-domain data and financial-domain data into one. Notably, the distribution of data in hybrid-tuning is non-conventional: financial DAP data comprises only a small portion of 13%. This prompts a pertinent question in line with the investigation on efficient DAP in (xie2023efficient, ): Is a large DAP dataset necessary for developing a domain-specific LLM?
Scientific Domain. Vertical scientific LLMs (Taylor2022Galactica, ; Yin2023FORGE, ; Xie2023DARWIN, ; Zhang2024SciGLM, ) span many subjects, including astronomy (Nguyen2023AstroLLaMA, ; Perkowski2024AstroLLaMAChat, ), mathematics (Azerbayev2023LLEMMA, ; Yu2023Outcome, ; luo2023wizardmath, ; yue2023mammoth, ; Gou2023ToRA, ), geology (Gao2023GLLAVA, ; roberts2023gpt4geo, ; lin2023geogalactica, ; wang2023nearrealtime, ), chemistry and physics (Rubungo2023LLM-Prop, ), biology (Yang2023PLLaMa, ; Zhao2023GIMLET, ; Cao2023InstructMol, ; Abdine2023Prot2Text, ; xTrimoPGLM2024Chen, ; Bi2023OCEANGPT, ; Zheng2023MarineGPT, ). However, among all the studies listed above, only a small fraction of them adopt the technique of DAP.
@@
OceanGPT (Bi2023OCEANGPT, ) is the first LLM tailored specifically for the ocean domain. It performs DAP on a raw corpus of ocean science literature, prioritizing recent research and historically significant works. K2 (deng2023learning, ) pioneers the development of a foundational language model tailored specifically for geoscience. It aggregates geoscience open access literature and Earth science-related Wikipedia pages for DAP. Following this, it undergoes multi-task instruction tuning utilizing LoRA (hu2022lora, ) on both a general instruction tuning dataset and the GeoSignal benchmark introduced within the K2 framework. AstroLlama (Nguyen2023AstroLLaMA, ) gathers abstracts solely from astronomy papers on arXiv and proceeds pre-training. It observes an improved perplexity on the domain of scholarly astronomy, without providing more quantitative evaluation. MarineGPT (Zheng2023MarineGPT, ) is a multi-modal LLM designed specifically for the marine domain. During DAP, MarineGPT incorporates 5 million marine image-text pairs to imbue domain knowledge. This involves training the parameters of a Q-Former (li2023blip2, ) between the frozen visual encoder (dosovitskiy2020image, ) and text decoder (touvron2023llama, ).
@@
Another branch of methods proactively integrate in the replay of the general-domain data to mitigate vertical forgetting. GeoGalactica (lin2023geogalactica, ) introduces a series of LLMs tailored for geoscience. In the DAP phase, besides the 52-billion-token geoscience corpus, Arxiv papers and Codedata are incorporated, with a mixing ratio of 8:1:1. The authors believe that the inclusion of the Codedata during the model’s pre-training can significantly boost the reasoning ability of the LLMs. Although GeoGalactica pinpoints challenges of DAP, including overfitting, catastrophic forgetting, maintaining the training stability, and convergence speed, it does not further provide empirical evidence supporting the inclusion of the Codedata, or deploying specific measures to address the challenges proposed above. Llemma (Azerbayev2023LLEMMA, ) focuses on mathematics, initialized from Code Llama (rozière2024code, ), and undergoes DAP on a blend of the 55 billion mathematical pre-training dataset and general domain data at the ratio of 19:1. In contrast, PLlama (Yang2023PLLaMa, ), designed for plant science, mixes domain-specific and general-domain data at the ratio of 9:1.
@@
Code Domain. The development of LLMs for automatic code filling, debugging, and generation holds significant practical importance (moradidakhel2023github, ; sun2024survey, ). These advancements cover various frameworks, including encoder-only (moradidakhel2023github, ), encoder-decoder (wang2021codet5, ; wang2023codet5plus, ; chai2023erniecode, ), and decoder-only (nijkamp2022codegen, ; nijkamp2023codegen2, ; zheng2023codegeex, ; chen2021evaluating, ; li2023starcoder, ; lozhkov2024starcoder, ; guo2024deepseekcoder, ). In the era of LLMs, there’s a growing trend towards decoder-only architectures (sun2024survey, ), leveraging models pre-trained on general natural language like Llama (touvron2023llama, ; touvron2023llama2, ). Consequently, there’s a shift in the training objective from utilizing code structures to simpler tasks like next token prediction and infilling.
@@
From the perspective of CL, the code domain presents unique advantages and challenges for DAP, compared to other domains discussed so far. On one hand, its hierarchical structure (general domain corpus $\rightarrow$ multi-language code $\rightarrow$ specific programming language) provides an ideal training pipeline for DAPs (rozière2024code, ), offering potential for more efficient training strategies. On the other hand, programming languages adhere to strict grammars, unlike the fuzzy and context-dependent natural language. Consequently, language models should ideally leverage these structures through tailored designs, and adopting the same training objectives as for natural languages may yield sub-optimal results. Therefore, many existing studies omit DAP (wang2021codet5, ; wang2023codet5plus, ; luo2023wizardcoder, ; muennighoff2024octopack, ; jiang2023selfevolve, ; wei2023magicoder, ; zhuo2024astraios, ; di2023codefuse, ; li2024instructcoder, ). In the following section, we will introduce existing code LLMs that employ DAP before the final downstream tasks, discussing both their common attributes and unique characteristics.
@@
Representing a series of notable works that focus solely on adaptation to target domains, CodeGen (nijkamp2022codegen, ) comprises a suite of LLMs designed for natural language (CodeGen-NL), multi-lingual programming languages (CodeGen-Multi), and mono-lingual programming languages (CodeGen-Mono). These models are trained sequentially, with each subsequent model initialized from the previous one trained on more general-domain data. Comment-Aug (song2024code, ) addresses the challenge of aligning programming languages with natural languages (PL-NL alignment) by performing DAP on the code augmented with generated additional comments. StarCoder (li2023starcoder, ) introduces two models: StarCoderBase and StarCoder. StarCoderBase is initially trained on a mixed dataset comprising various programming languages without significant reweighting on the data. Subsequently, StarCoderBase undergoes further fine-tuning on an additional 35 billion tokens of Python code, resulting in the development of StarCoder. DeepSeek-Coder-v1.5 (guo2024deepseekcoder, ) originates from DeepSeek-LLM (deepseekai2024deepseek, ) and undergoes pre-training on 2 trillion tokens, comprising 87% source code, 10% English code-related natural language, and 3% Chinese natural language corpus. Initialization from a general-domain LLM results in improved performance across various tasks, including natural language and mathematical reasoning, with minimal performance degradation on coding tasks, which underscores the efficacy of DAP.